OpenStudy (anonymous):

Use the properties of logarithms to express as a sum, difference, and or mulitple of logarithms. log(x(squre root of x-3)

OpenStudy (anonymous):

logx + (1/2)log(x-3)

OpenStudy (anonymous):

\[\log(A*B)=\log(A)+\log(B)\]

OpenStudy (anonymous):

\[\log(x^a)=a*\log(x)\]

OpenStudy (anonymous):

so log(xsqrt(x-3))= log(x)+log(sqrt(x-3)= log(x)+log((x-3)^(1/2))= log(x)+(1/2)log(x-3)